How do you write an inequality of line AB that has point (4,8) and (-10,1) in standard form with integers?
Answer by stanbon(75887)

How do you write an inequality of line AB that has point (4,8) and (-10,1) in standard form with integers?
---
1st: Write an equality; it is the boundary between inequalities.
---
slope = (8-1)/(4--10) = 7/14 = 1/2
---
Form: y = mx+b
8 = (1/2)(4) + b
8 = 2 + b
b = 6
-----
Equation: y = (1/2)x+6
---
Standard Form:
2y = x + 12
x - 2y = -12
------
Corresponding inequalities:
x-2y<-12
OR
x-2y>-12
===================
Cheers,
Stan H.